OpenStudy (anonymous):

Find the axis of symmetry for: 2x^2-4x+2 Isn't it something like: x=1?

OpenStudy (aum):

This is a vertical parabola that opens upward. So the axis of symmetry will be a vertical line of the form x = constant that passes through the vertex of the parabola. What is the x-coordinate of the vertex of the parabola?

OpenStudy (campbell_st):

for an equation \[y = ax^2 + bx + c\] the axis of symmetry is \[b = \frac{-b}{2a}\] you have a = 2 and b = -4 so substitute and find it

OpenStudy (campbell_st):

oops should read \[x = \frac{-b}{2a}\]

OpenStudy (anonymous):

For the vertex I got (1,0)

OpenStudy (anonymous):

That's why I said x=1

OpenStudy (campbell_st):

well that's another valid method... putting the equation in vertex form y=2(x - 1)^2 there are lots of methods... both show the line of symmetry is x = 1